# Every finite set of natural numbers is realizable as algebraic periods of a Morse–Smale diffeomorphism

### v1

@@ Abstract.
A given self-map $f\colon M\to M$ of a compact manifold determines the sequence $\{L(f^{n})\}_{n=1}^{\infty}$, of the Lefschetz numbers of its iterations. We consider its dual sequence $\{a_{n}(f)\}_{n=1}^{\infty}$ given by the Möbius inversion formula. The set ${\mathcal{A}P}(f)=\{n\in\mathbb{N}\colon a_{n}(f)\neq 0\}$ is called the set of algebraic periods.
We solve an open problem existing in literature by showing that for every finite subset ${\mathcal{A}}$ of natural numbers there exist an orientable surface $S_{\rm g}$, as well as a non-orientable surface $N_{\mathrm{g}}$, of genus ${\rm g}$, and a Morse–Smale diffeomorphism $f$ of this surface such that $\mathcal{AP}(f)=\mathcal{A}$. For such a map it implies the existence of points of a minimal period $n$ for each odd $n\in\mathcal{A}$. For the orientation-reversing Morse–Smale diffeomorphisms of $S_{\mathrm{g}}$ we identify strong restrictions on ${\mathcal{A}P}(f)$. Our method also provides an estimate of the number of homotopy classes of Morse–Smale diffeomorphisms, which is exponential in $\mathrm{g}$.
@@ 1. Introduction
Let $f:X\to X$ be a self-map of a topological space $X$. One of the classical problems of the topological theory of dynamical systems is to find $n$-periodic points of $f$, i.e. $x\in X$ such that $f^{n}(x)=x$ for some $n\in\mathbb{N}$. If $f^{k}(x)\neq x$ for $k<n$, then $n$ is called the minimal period of $f$ at $x$, and $x$ is a point of period $n$. Then the set $\{x,f(x),f^{2}(x),\,\dots\,,f^{n-1}(x)\}$ is called the orbit of $x$. We denote by $P^{n}(f)={\rm Fix}(f^{n})$ the set of all $n$-periodic points, and by $P_{n}(f)$ the set of all points of period $n$. We have $P_{n}(f)=P^{n}(f)\setminus{\underset{k|n,k<n}{\,\bigcup\,}}P^{k}(f)$.
@@
One of the natural problems in the field of periodic point theory is the following question: suppose that ${\mathcal{A}}$ is a subset of natural numbers, is it possible to realize it as the minimal periods of a self-map of some topological space $X$? This problem has been extensively studied (mainly from the arithmetical point of view) by many authors (cf. [EPPW] and the expository paper [BGW]) even in the stronger version, when the number of orbits of each minimal period is also given. In general, it turns out that there are no obstacles for such realizability for very general class of maps and relatively simple spaces. In particular any (also infinite) sequence of natural numbers can be realized for smooth maps on two-dimensional torus $\mathbb{T}^{2}$ [Winds]. However, the approach in [Winds] provides the realization in the homotopy class of identity map, so the periodic points are not detectable by topological/homotopical methods. In this paper we address a more subtle version of the problem from the perspective of homotopy invariants and for a more restricted class of maps (see Main Problem below). We also discuss the relations of our problem with the existence of periodic points (refer to Proposition 3.8 and Corollary 3.9).
1.1Main Problem.
@@
Remind that every Morse–Smale diffeomorphism has only a finite-number of periodic points by the definition (cf. [Smale]).
Below, we sketch the motivation to study the problem in such a formulation. The tools for posing and solving this problem are closely related to the classical concept of algebraic topology, namely the Lefschetz number $L(f)$, which is homotopy invariant (see [Je-Ma] for a comprehensive overview and details).
With the sequence $\{L(f^{n})\}_{n=1}^{\infty}$, we associate the dual sequence given by the Möbius inversion formula
@@
where $\mu:\mathbb{N}\to\{-1,0,1\}$ is the Möbius function. From the Möbius inversion formula, it follows that $\;\;L(f^{n})=\sum_{k\mid n}\,k\,a_{k}(f).$ By the definition, $a_{n}(f)\in\mathbb{Q}$, but in fact, they are integers due to the Dold’s theorem (cf. [Do]).
@@
Let $X$ be a finite set and $f\colon X\to X$ a map. Then $L(f^{n})=|P^{n}(f)|$, and consequently $a_{n}(f)=|P_{n}(f)|$. Thus, in this example $a_{n}(f)\neq 0$ implies $P_{n}(f)\neq\emptyset$ which gives $n\in{\rm Per}(f)$.
@@
In general, non-vanishing of $a_{n}(f)$ does not imply the existence of points of period $n$. However, for some special classes of smooth self-maps of manifolds, such as the Morse–Smale diffeomorphisms, or transversal $C^{1}$ maps, $a_{n}(f)\neq 0$ implies $P_{n}(f)\neq\emptyset$ if $n$ is odd, or $P_{n}(f)\cup P_{\frac{n}{2}}(f)\neq\emptyset$ if $n$ is even (see Proposition 3.8). The above led Llibre, Sirvent and other collaborators to study ${\mathcal{A}P}(f)$, or more precisely its odd part ${\mathcal{A}P}_{odd}(f)={\mathcal{A}P}(f)\cap\{2\mathbb{N}-1\}$. They had been considering situations when it is computable by direct algebraic topology methods (see [Ll-Si3] for an exposition, and the bibliography of the results of this group up to 2013). In fact, instead of algebraic periods, they considered a subset of natural numbers called ”the set of minimal Lefschetz periods” denoted by ${\rm MPer}_{L}(f)$ (see Definition 5.1). Later, these authors noted that this set does not contain even numbers (cf. [Ll-Si2]). Recently Graff, Lebiedź and Myszkowski showed in [GLM] that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$, which justifies to use only the notion of algebraic periods. The proof of [GLM] uses the $k$-periodic expansion (cf. the formula (2)) and is based on the local expression of the fixed point index as a combination of basic periodic sequences. As an appendix, we provide a purely algebraic proof of the fact that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$ in Theorem 5.2.
@@
A direct inspiration for this work was the papers [Ll-Si1, Ll-Si3] and the following question posed there:
@@
Can any finite set ${\mathcal{A}}$ of odd positive integers be the minimal set of Lefschetz periods for a $C^{1}$ Morse–Smale diffeomorphism on some orientable/non-orientable compact surface without boundary with a convenient genus $\mathrm{g}$.
@@
As stated by the authors in [Ll-Si1], in all their works they do not characterize the sets of homotopy classes of Morse–Smale diffeomorphisms. They only described a set called the minimal sets of Lefschetz periods ${\rm MPer}_{L}(f)$ (Definition 5.1), i.e. ${\mathcal{A}P}(f)\setminus 2\mathbb{N}$ of quasi-unipotent homeomorphisms (i.e. homeomorphisms for which the induced linear map $H_{*}(f)$ on the homology of $M$ has spectrum consisting of the roots of unity only). Remind that if $f:M\to M$ is a Morse–Smale diffeomorphism, then $f$ is quasi-unipotent by the Shub result [Shub], but not conversely. Consequently their results gave only the necessary conditions, i.e. the restrictions on the algebraic periods of Morse–Smale diffeomorphisms never providing a geometrical realization. These restrictions were obtained by purely algebraic computation of the Lefschetz zeta function based on known forms of the homology groups, or cohomology rings, of studied manifolds.
@@
In this work, by a direct geometrical argument we give a complete positive answer to the Problem 1.1, i.e. a stronger version of the stated above question Libre and Sirvent, proving that for any finite set ${\mathcal{A}}\subset\mathbb{N}$ there exists an orientable surface of genus $\mathrm{g}$ and a preserving orientation diffeomorphism $f:S_{\mathrm{g}}\to S_{\mathrm{g}}$ such that ${\mathcal{A}P}(f)={\mathcal{A}}$ (Theorem 1.4). In the case we demand that the realization is performed in the class of reversing orientation Morse–Smale diffeomorphism we find strong restrictions on ${\mathcal{A}P}(f)$, namely ${\mathcal{A}P}(f)\cap\{2\mathbb{N}-1\}=\emptyset$, i.e. the set of algebraic periods consists only of even numbers if $f$ is a reversing orientation diffeomorphism of an orientable surface, which was first shown by A. Myszkowski in his unpublished Ph. D. thesis [Mysz]. Here we provide another proof of this fact. This result corrects the wrong statement regarding $S_{3}$ in [Ll-Si1] (Theorem 5.7). Finally, we show that for any finite set ${\mathcal{A}}\subset\mathbb{N}$ there exists a non-orientable surface of genus $\mathrm{g}$ and a Morse–Smale diffeomorphism $f:N_{\mathrm{g}}\to N_{\mathrm{g}}$ such that ${\mathcal{A}P}(f)={\mathcal{A}}$ (Theorem 1.4).
@@
Let ${\mathcal{A}}$ be a finite set of natural numbers. There exist a Morse–Smale diffeomorphism $f$ on a closed surface such that ${\mathcal{A}}={\mathcal{A}P}(f)$. The surface can be chosen both orientable and non-orientable. Moreover, $f$ can be orientation-reversing (in orientable case) if and only if ${\mathcal{A}}\subset 2{\mathbb{N}}$.
One of the consequences of the construction provided in Theorem 1.4 is an exponential growth in $\mathrm{g}$ of the number of different algebraically finite type mapping classes, so consequently of homotopy classes of the Morse–Smale diffeomorphisms on a surface of a given genus (Theorem 4.7).
Another important statement that we draw from our main theorem is the fact that for Morse–Smale diffeomorphisms (as well as more for a more general class of transversal maps) we are able to provide not only the realization of algebraic periods $n\in{\mathcal{A}P}(f)$ but also a $n$-periodic point for $n$ odd, see Subsection 3.1.
@@
In this subsection, we give a brief survey of known results that describe the homotopy classes of homeomorphisms, thus diffeomorphisms, of surfaces which contain the Morse–Smale diffeomorphisms. From now on, we assume that our surface is hyperbolic, i.e. it is of genus $\mathrm{g}\geq 2$. The case of surfaces with non-negative Euler characteristic was studied separately in [da Rocha].
As the first we remind the Thurston classification theorem [FLP, Farb-Marg], called also the Nielsen–Thurston classification theorem.
Let $f$ be a homeomorphism of a closed orientable surface $S_{\mathrm{g}}$ of genus $\mathrm{g}\geq 2$. Denote by $[f]$ the set of all homeomorphisms of the surface $S_{\mathrm{g}}$ that are homotopic to $f$ (the homotopy class of homeomorphisms containing $f$). According to the Nielsen–Thurston classification (see, [FLP, Thms 11.6,   11.7] and [Gri-Mor-Poch]) the set of all homotopy classes of homeomorphisms on $S_{\mathrm{g}}$ is represented as the union of four disjoint subsets T1, T2, T3, T4 distinguished by the conditions described below.
@@
    if $[f]\in T2$, then $[f]$ contains a reducible non-periodic homeomorphism of algebraically finite order;
@@
    if $[f]\in T3$, then $[f]$ contains a reducible homeomorphism which is neither periodic nor of algebraically finite order;
@@
The classes $T1$ and $T2$ are called algebraically finite type in the original Nielsen terminology. The Nielsen–Thurston theory was developed for orientable surfaces of genus $\mathrm{g}\geq 2$. The set of homotopy classes of Morse–Smale diffeomorphisms of the torus, and also of non-orientable surfaces of genus $\leq 1$, i.e. the projective space and Klein bottle is described in [da Rocha] (see Theorem 2.3).
Roughly speaking, in the reducible case (T2 and T3), Theorem 2.1 states that one can cut a surface along the invariant collection of closed curves obtaining a “canonical form”. This constitutes a collection of surfaces with boundary, satisfying that for each of them some iteration of the map is its self-homeomorphism. As a consequence, any mapping class can be reduced into either only periodic pieces (T2) or periodic and pseudo-Anosov pieces (T3).
The analogue of Theorem 2.1 for non-orientable surfaces was shown by Yingqing Wu in [WuYing] (see [Par] for an exhaustive exposition). The statement of the mentioned version has the same formulation and the main idea of its proof is to proceed with $\tilde{f}^{+}:S_{\mathrm{g}}\to S_{\mathrm{g}}$ the preserving orientation lift of a homeomorphism $f:N_{\mathrm{g}}\to N_{\mathrm{g}}$, where $p:S_{\mathrm{g}}\to N_{\mathrm{g}}$ is two-sheets orientable cover of the non-orientable surface $N_{\mathrm{g}}$.
Before next consideration, let us remind the Baer–Epstein theorem (cf. [Farb-Marg]):
@@
For homeomorphisms of a surface the homotopy is equivalent to the isotopy.
@@
In the forthcoming part of the paper, we will base essentially on the following result of Luis F. da Rocha [da Rocha, THM A]
@@
If $M^{2}$ is a two dimensional compact connected boundaryless manifold (orientable or not) with a negative Euler characteristic, then the condition that $f$ is of algebraically finite type is necessary and sufficient for an isotopy class of $f$ of $M^{2}$ to be Morse–Smale i.e. contains a Morse–Smale representant).
In the case $\chi(M^{2})\geq 0$ we have the following situation:
@@
    if $M^{2}$ is the two dimensional torus, an isotopy class is Morse–Smale if and only if:
@@
Theorem 2.3 was reproved by A.N. Bezdenezhykh and V.Z. Grines [Bez-Grin] for the class $T1$, and by V. Grines, A. Morozov, and O. Pochinka [Gri-Mor-Poch] for the class $T2$ of homeomorphisms of orientable surfaces. The authors of the referred articles used subtle analytical methods, and providing constructions of such Morse–Smale diffeomorphisms solving more advanced questions.
Fact 2.5.
For a homeomorphism of $f\colon M\to M$ of a compact surface $M$ the following are equivalent:
  1. (1)
    $\{L(f^{n})\}$ is bounded,
  2. (2)
@@
The equivalence between the first and the second condition is given in Theorem [Je-Ma, 3.1.46] (cf. [Bab-Bog]). Since the Lefschetz numbers of iterations of a diffeomorphism $f$ of a surface are equal to $1+\mathrm{t}rH_{1}(f)^{n}+\pm 1$ in the orientable case, and respectively $1+\mathrm{t}rH_{1}(f)^{n}$ in the non-orientable case, the first and third conditions are equivalent. ∎
However, let us mention here that the spectral radius of the representation $f\mapsto H_{1}(f)$ is not a strong enough invariant to determine whether $[f]\in T1\cup T2$ (cf. [FLP] or [Farb-Marg] for the Thurston example of a pseudo-Anosov homeomorphism $f$ with $H_{1}(f)={\rm id}$.

### v2

@@ Abstract.
A given self-map $f\colon M\to M$ of a compact manifold determines the sequence $(L(f^{n}))$ of the Lefschetz numbers of its iterations. We consider its dual sequence $(a_{n}(f))$ given by the Möbius inversion formula. The set ${\mathcal{A}P}(f)=\{n\in\mathbb{N}\colon a_{n}(f)\neq 0\}$ is called the set of algebraic periods.
We solve an open problem existing in literature by showing that for every finite subset ${\mathcal{A}}$ of natural numbers, there exist an orientable surface $S_{\rm g}$, as well as a non-orientable surface $N_{\mathrm{g}}$, of genus ${\rm g}$, and a Morse–Smale diffeomorphism $f$ of this surface such that $\mathcal{AP}(f)=\mathcal{A}$. For such a map it implies the existence of points of a minimal period $n$ for each odd $n\in\mathcal{A}$. For the orientation-reversing Morse–Smale diffeomorphisms of $S_{\mathrm{g}}$, we identify strong restrictions on ${\mathcal{A}P}(f)$. Our method also provides an estimate of the number of conjugacy classes of mapping classes containing Morse–Smale diffeomorphisms, which is exponential in $\mathrm{g}$.
@@ 1. Introduction
Let $f\colon X\to X$ be a self-map of a topological space $X$. One of the classical problems of the topological theory of dynamical systems is to find $n$-periodic points of $f$, i.e. $x\in X$ such that $f^{n}(x)=x$ for some $n\in{\mathbb{N}}$. If $f^{k}(x)\neq x$ for $k<n$, then $n$ is called the minimal period of $f$ at $x$, and $x$ is a point of period $n$. Then, the set $\{x,f(x),f^{2}(x),\,\dots\,,f^{n-1}(x)\}$ is called the orbit of $x$. We denote by $P^{n}(f)={\rm Fix}(f^{n})$ the set of all $n$-periodic points, and by $P_{n}(f)$ the set of all points of period $n$. We have $P_{n}(f)=P^{n}(f)\setminus{\underset{k|n,k<n}{\,\bigcup\,}}P^{k}(f)$.
@@
One of the natural problems in the field of periodic point theory is the following question: Supposing that ${\mathcal{A}}$ is a subset of natural numbers, is it possible to realize it as the minimal periods of a self-map of some topological space $X$? This problem has been extensively studied (mainly from the arithmetical point of view) by many authors (cf. [6] and the expository paper [4]), even in the stronger version when the number of orbits of each minimal period is also given. In general, it turns out that there are no obstacles for such realizability for a very general class of maps and relatively simple spaces. In particular, any (also infinite) sequence of natural numbers can be realized for smooth maps on a two-dimensional torus $\mathbb{T}^{2}$ [29]. However, the approach in [29] provides the realization in the homotopy class of the identity map, so the periodic points are not detectable by topological/homotopical methods. In this paper, we address a more subtle version of the problem from the perspective of homotopy invariants and for a more restricted class of maps (see Main Problem below). We also discuss the relations of our problem with the existence of periodic points (refer to Proposition 3.8 and Corollary 3.9).
Main Problem.
@@
Recall that every Morse–Smale diffeomorphism has only a finite number of periodic points, which follows easily from its definition (cf. [27]).
Below, we sketch the motivation to study the problem in such a formulation. The tools for posing and solving this problem are closely related to the classical concept of algebraic topology, namely the Lefschetz number $L(f)$, which is homotopy invariant (see [14] for a comprehensive overview and details).
With the sequence $(L(f^{n}))$, we associate the dual sequence $(a_{n}(f))$ given by the Möbius inversion formula
@@
where $\mu\colon{\mathbb{N}}\to\{-1,0,1\}$ is the Möbius function. From the Möbius inversion formula, it follows that $\;\;L(f^{n})=\sum_{k\mid n}\,k\,a_{k}(f).$ By definition, $a_{n}(f)\in{\mathbb{Q}}$, but in fact they are integers due to Dold’s theorem (cf. [5]).
@@
Let $X$ be a finite set and $f\colon X\to X$ a map. Then, $L(f^{n})=|P^{n}(f)|$, and consequently $a_{n}(f)=|P_{n}(f)|$. Thus, in this example, $a_{n}(f)\neq 0$ implies $P_{n}(f)\neq\emptyset$ which gives $n\in{\rm Per}(f)$.
@@
In general, the non-vanishing of $a_{n}(f)$ does not imply the existence of points of period $n$. However, for some special classes of smooth self-maps of manifolds, such as the Morse–Smale diffeomorphisms or transversal $C^{1}$ maps, $a_{n}(f)\neq 0$ implies $P_{n}(f)\neq\emptyset$ if $n$ is odd, or $P_{n}(f)\cup P_{\frac{n}{2}}(f)\neq\emptyset$ if $n$ is even (see Proposition 3.8). The above led Llibre, Sirvent, and other collaborators to study ${\mathcal{A}P}(f)$, or more precisely its odd part ${\mathcal{A}P}_{odd}(f)={\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)$. They had been considering situations when it is computable by direct algebraic topology methods (see [18] for an exposition, and the bibliography of the results of this group up to 2013). In fact, instead of algebraic periods, they considered a subset of natural numbers called “the minimal set of Lefschetz periods” denoted by ${\rm MPer}_{L}(f)$ (see Definition A.1). Later, these authors noted that this set does not contain even numbers (cf. [17]). Recently, Graff, Lebiedź, and Myszkowski showed in [10] that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$, which justifies to use only the notion of algebraic periods. The proof of [10] uses the periodic expansion (cf. the formula (2)) and is based on the local expression of the fixed point index as a combination of basic periodic sequences. In the appendix, we provide a purely algebraic proof of the fact that ${\rm MPer}_{L}(f)={\mathcal{A}P}_{odd}(f)$ in Theorem A.2.
@@
A direct inspiration for this work was the papers [16, 18] and the following question posed there:
@@
Can any finite set ${\mathcal{A}}$ of odd positive integers be the minimal set of Lefschetz periods for a $C^{1}$ Morse–Smale diffeomorphism on some orientable/non-orientable compact surface without boundary with a convenient genus $\mathrm{g}$?
@@
As stated by the authors in [16], in all their works they do not characterize the sets of homotopy classes of Morse–Smale diffeomorphisms. They only described the sets called the minimal sets of Lefschetz periods ${\rm MPer}_{L}(f)$ (Definition A.1), i.e. ${\mathcal{A}P}(f)\setminus 2{\mathbb{N}}$ of quasi-unipotent homeomorphisms (i.e. homeomorphisms for which the induced linear map $H_{*}(f)$ on the homology of $M$ has spectrum consisting of the roots of unity only). Remind that if $f\colon M\to M$ is a Morse–Smale diffeomorphism, then $f$ is quasi-unipotent by the Shub result [25], but not conversely. Consequently, their results gave only the necessary conditions, i.e. the restrictions on the algebraic periods of Morse–Smale diffeomorphisms never providing a geometrical realization. These restrictions were obtained by purely algebraic computation of the Lefschetz zeta function based on known forms of the homology groups, or cohomology rings, of studied manifolds.
@@
In this work, by a direct geometrical argument we give a complete positive answer to the Main Problem, i.e. a stronger version of the stated above question Llibre and Sirvent, proving that for any finite set ${\mathcal{A}}\subset{\mathbb{N}}$ there exists an orientable surface of genus $\mathrm{g}$ and a preserving orientation diffeomorphism $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ such that ${\mathcal{A}P}(f)={\mathcal{A}}$ (Theorem 1.3). In the case in which demand that the realization is performed in the class of reversing orientation Morse–Smale diffeomorphism, we find strong restrictions on ${\mathcal{A}P}(f)$, namely ${\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)=\emptyset$, i.e. the set of algebraic periods consists only of even numbers if $f$ is a reversing orientation diffeomorphism of an orientable surface, which was shown by A. Myszkowski in his unpublished PhD thesis [20], but the result follows also from the paper of Blanchard and Franks [3], and we present this approach in Theorem 3.2. This result corrects the wrong statement regarding $S_{3}$ in [16] (Theorem 5.7). Finally, we show that for any finite set ${\mathcal{A}}\subset{\mathbb{N}}$, there exists a non-orientable surface of genus $\mathrm{g}$ and a Morse–Smale diffeomorphism $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ such that ${\mathcal{A}P}(f)={\mathcal{A}}$ (Theorem 1.3).
@@
Let ${\mathcal{A}}$ be a finite set of natural numbers. There exist a Morse–Smale diffeomorphism $f$ on a closed surface such that ${\mathcal{A}}={\mathcal{A}P}(f)$. The surface can be chosen both orientable and non-orientable. Moreover, $f$ can be orientation-reversing (in the orientable case) if and only if ${\mathcal{A}}\subset 2{\mathbb{N}}$.
One of the consequences of the construction provided in Theorem 1.3 is an exponential growth in $\mathrm{g}$ of the number of different conjugacy classes of algebraically finite type mapping classes, so, consequently, of conjugacy classes of homotopy classes of the Morse–Smale diffeomorphisms on a surface of a given genus (Theorem 4.7).
Another important statement that we draw from our main theorem is the fact that for Morse–Smale diffeomorphisms (as well as more for a more general class of transversal maps) we are able to provide not only the realization of algebraic periods $n\in{\mathcal{A}P}(f)$, but also an $n$-periodic point for $n$ odd, see Subsection 3.1.
@@
In this section, we give a brief survey of known results that describe the homotopy classes of homeomorphisms, thus diffeomorphisms, of surfaces which contain the Morse–Smale diffeomorphisms. From now on, we assume that our surface is hyperbolic, i.e. it is of genus $\mathrm{g}\geq 2$. The case of surfaces with non-negative Euler characteristic was studied separately in [23].
First, we recall the Thurston classification theorem [7, 8], also known as the Nielsen–Thurston classification.
Let $f$ be a homeomorphism of a closed orientable surface $S_{\mathrm{g}}$ of genus $\mathrm{g}\geq 2$. Denote by $[f]$ the set of all homeomorphisms of the surface $S_{\mathrm{g}}$ that are homotopic to $f$ (the homotopy class of homeomorphisms containing $f$). According to the Nielsen–Thurston classification (see [8, Thms 11.6, 11.7] and [13]), the set of all homotopy classes of homeomorphisms on $S_{\mathrm{g}}$ is represented as the union of four disjoint subsets T1, T2, T3, and T4 distinguished by the conditions described below.
@@
    if $[f]\in T2$, then $[f]$ contains a reducible non-periodic homeomorphism of algebraically finite type;
@@
    if $[f]\in T3$, then $[f]$ contains a reducible homeomorphism which is neither periodic nor of algebraically finite type;
@@
The classes $T1$ and $T2$ are called algebraically finite type in the original Nielsen terminology. Nielsen–Thurston theory was developed for orientable surfaces of genus $\mathrm{g}\geq 2$. The case $g=0$ (the sphere) is trivial, and for the case $g=1$ (the torus) the classification of homotopy classes of its homeomorphisms is provided algebraically by elements of $SL(2,{\mathbb{Z}})$, which is a classical fact (cf. [7]).
Roughly speaking, in the reducible case (T2 and T3), Theorem 2.1 states that one can cut a surface along the invariant collection of closed curves, obtaining a “canonical form”. This constitutes a collection of surfaces with boundary, satisfying that for each of them some iteration of the map is its self-homeomorphism. As a consequence, any such mapping class can be reduced into either only periodic pieces (T2) or periodic and pseudo-Anosov pieces (T3). Thus, a mapping class is of algebraically finite type if all its pieces (even if one) in its Nielsen–Thurston reduction are periodic.
The analogue of Theorem 2.1 for non-orientable surfaces was shown by Yingqing Wu in [30] (see [22] for an exhaustive exposition). The statement of the mentioned version has the same formulation, and the main idea of its proof is to proceed with $\tilde{f}^{+}\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ the preserving orientation lift of a homeomorphism $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$, where $p\colon S_{\mathrm{g}}\to N_{\mathrm{g}}$ is a two-sheets orientable cover of the non-orientable surface $N_{\mathrm{g}}$.
Before our next consideration, let us recall the Baer–Epstein theorem (cf. [7]).
@@
For homeomorphisms of a surface, the homotopy is equivalent to the isotopy.
@@
The forthcoming part of the paper is based on the following result of Luis F. da Rocha [23, THM A]
@@
If $M^{2}$ is a two-dimensional compact connected boundaryless manifold (orientable or not) with a negative Euler characteristic, then the condition that $f$ is of algebraically finite type is necessary and sufficient for an isotopy class of $f$ of $M^{2}$ to be Morse–Smale, i.e. it contains a Morse–Smale representant.
In the case $\chi(M^{2})\geq 0$, we have the following situation:
@@
    if $M^{2}$ is the two dimensional torus, an isotopy class is Morse–Smale if and only if
@@
Theorem 2.3 was reproved by A.N. Bezdenezhykh and V.Z. Grines [2] for the class $T1$, and by V. Grines, A. Morozov, and O. Pochinka [13] for the class $T2$ of homeomorphisms of orientable surfaces. The authors of the referred articles used subtle analytical methods, and provided constructions of such Morse–Smale diffeomorphisms, solving more advanced questions.
Proposition 2.5.
For a homeomorphism of $f\colon M\to M$ of a compact surface $M$, the following are equivalent:
  • 1)
    $(L(f^{n}))$ is bounded,
  • 2)
@@
The equivalence between the first and the second condition is given in [14, Theorem 3.1.46] (cf. [1]). Since the Lefschetz numbers of iterations of a diffeomorphism $f$ of a surface are equal to $1-\mathrm{t}rH_{1}(f)^{n}+\pm 1$ in the orientable case, and respectively $1-\mathrm{t}rH_{1}(f)^{n}$ in the non-orientable case, the first and third conditions are equivalent. ∎
However, let us mention here that the spectral radius of the representation $f\mapsto H_{1}(f)$ is not a strong enough invariant to determine whether $[f]\in T1\cup T2$ (cf. [8] or [7] for the Thurston example of a pseudo-Anosov homeomorphism $f$ with $H_{1}(f)={\rm id}$).